OpenStudy (anonymous):

You randomly select a month of the year. What is the probability that you select a month with 28 or 30 days in it ?

OpenStudy (anonymous):

How many months have 28 days? How many months have 30 days? Also, do you mean months that ONLY have 28 days?

OpenStudy (anonymous):

only one month has 28 days . then the rest has 3 days . Well not sure the question states 28 and 30 days in it

OpenStudy (anonymous):

The reason I ask is because all months have 28 days (not just February). However, if the question means how many have ONLY 28 days, then Feb is the only one. Agreed?

OpenStudy (anonymous):

Yes , but not even myself understands the questions !

OpenStudy (anonymous):

Let's assume the question takes on the ONLY form. There are 12 months in total. Only February has ONLY 28 days. (one month) April, June, September, and November have ONLY 30 days. (4 months) Say \(A\) is the event that a month has ONLY 28 days, and \(B\) is the event that a month has ONLY 30 days. Then \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] \(P(A\cap B)=0\) because you can't have a month that has ONLY 28 days and ONLY 30 days; these events are mutually exclusive. \[P(A\cup B)=\frac{1}{12}+\frac{4}{12}-0=\frac{5}{12}\]

OpenStudy (anonymous):

On the other hand, if the question is just talking about months that generally have 28 or 30 days, then the probability is 100% because all 12 months have at least 28 days.

OpenStudy (anonymous):

so the answer can vary ? depending on the significant of the question ? ugh still confused ! not sure what the meaning of the question really means

OpenStudy (anonymous):

It's a surprisingly tough question. I'd write down both answers :P

OpenStudy (anonymous):

Well sadly I don't have that option ! I could only choose one answer